OpenStudy (anonymous):

At a basketball game, a vender sold a combined total of 111 sodas and hot dogs. The number of sodas sold was two times the number of hot dogs sold. Find the number of sodas and the number of hot dogs sold.

rishavraj (rishavraj):

let total no of hot dogs sold = x so no of sodas sold = 2x 2x + x = 111 solve the equation

OpenStudy (anonymous):

x=37

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

yup...:)`

OpenStudy (anonymous):

is that it? @rishavraj

rishavraj (rishavraj):

no of hot dogs = 37 no of sodas = 2*37 = 74